a number plus 5 more than 3 times the number is 37 find the number
Answer by ikleyn(53751)

.

The problem says


    x + (3x+5) = 37.


Simplify and find the unknown number x


    4x + 5 = 37

    4x     = 37 - 5 = 32

     x              = 32/4 = 8.


ANSWER.  The number is 8.

Solved.